# Multi-Scale Data Assimilation in Turbulent Models

### v1

@@ VII Comparison with En4D-Var
Due to the chaotic nature of the shell model, any error - whether from the initial guess or from linearization - grows exponentially over time. As a result, the length of the assimilation window $T_{a}$ is typically limited to the eddy turnover time of the fastest measured shell. This poses a challenge for assimilating slower (i.e., lower-wavenumber) shells, which require longer windows. To address this, the 4D-Var process is performed sequentially: a long assimilation interval is divided into smaller windows, and background information is passed from one window to the next. A widely used strategy in numerical weather prediction to perform such sequential assimilation is the ensemble-based 4D-Var (En4D-Var) [25]. In this approach, multiple 4D-Var optimizations $\tilde{\bm{u}}^{(i,j)}(t)$ (with $j=1,\dots,L$ denoting the ensemble index) are run in parallel, using a prior ensemble $\hat{\bm{u}}^{(j)}(t)$ to provide the background-error covariance matrix $\hat{\bm{\Sigma}}_{\mathbb{C}}=\langle(\hat{\bm{u}}-\langle\hat{\bm{u}}\rangle_{L})(\hat{\bm{u}}-\langle\hat{\bm{u}}\rangle_{L})^{\dagger}\rangle_{L}$, and to propagate flow-dependent information across assimilation cycles. The modified cost function used in this ensemble-sequential version is:
$J^{(i,j)}_{n_{\text{seq}}}=$ $\left\|\hat{\bm{u}}^{(j)}(n_{\text{seq}}T_{a})-\tilde{\bm{u}}^{(i,j)}(n_{\text{seq}}T_{a})\right\|_{\hat{\bm{\Sigma}}_{\mathbb{C}}}$ (22)
$+\sum_{t=n_{\text{seq}}T_{a}}^{(n_{\text{seq}}+1)T_{a}}$ $\left\|\bm{z}(t)-\bm{H}_{c}\tilde{\bm{u}}^{(i,j)}(t)-\bm{v}^{(j)}(t)\right\|_{\bm{R}_{\mathbb{C}}}\,,$
@@
For the first sequence ($n_{\text{seq}}=0$), an ensemble $\hat{\bm{u}}^{(j)}$ is initialized as in Eq. (23) and 4D-Var calculations are solved for a maximum of 50 iterations or until the reduction in the cost function reaches a tolerance. For later sequences ($n_{seq}>0$), the initial guess is the same as the background information, which is available from the previous sequence as $\hat{\bm{u}}^{(j)}(n_{seq}T_{a})=\tilde{\bm{u}}^{(j)}(n_{seq}T_{a})$. We also note that the first sequence does not have background information provided.
@@
$\hat{u}_{n}^{(j)}(t_{0})=\begin{cases}\sqrt{\langle|u_{n}|^{2}\rangle_{T}}\left[\cos(\phi_{n}^{(j)})+i\sin(\phi_{n}^{(j)})\right]&n\neq m\\ u_{n}(t_{0})+\epsilon_{n}^{(j)}&n=m\end{cases}\ ,$ (23)
@@
Regarding the execution of the algorithm, a regularization term is added to improve numerical stability: $\left\|\tilde{\bm{u}}^{(i-1,j)}(0)-\tilde{\bm{u}}^{(i,j)}(0)\right\|_{\mathcal{P}}$, where $\mathcal{P}$ is a diagonal matrix defined as $\mathcal{P}(n,n)=N_{\text{reg}}$ for $n<15$ and $\mathcal{P}(n,n)=N_{\text{reg}}/(2E_{n})$ otherwise. This term does not significantly alter the cost function or its gradient (since $\tilde{\bm{u}}^{(i-1,j)}\approx\tilde{\bm{u}}^{(i,j)}$), but it helps to regularize the Hessian inversion.
@@
[…]
@@
In summary, ensemble-based methods are necessary to assimilate scales beyond those directly measured. Both EnKF and En4D-Var perform well when three consecutive shells are observed. However, En4D-Var is significantly more computationally expensive and less practical in operational settings, due to the difficulty of accessing or constructing the adjoint model.
@@ VIII Conclusions
Another promising direction involves integrating inflation optimization into a machine learning training framework. Since inflation effectively mimics the effect of larger ensemble sizes, learning the optimal inflation coefficients would allow us to simulate the behavior of massive ensembles without incurring their computational cost. A possible strategy is to adopt a solver-in-the-loop approach [46], where inflation parameters are updated through backpropagation throughout the assimilation time window, rather than in a single analysis step. This would enable a global optimization of the assimilation process and open new pathways for hybrid dynamical–ML data assimilation methods [47].
@@ Appendix A Runge-Kutta with exponential factor
$u_{n}(t+dt)=I_{n}^{2}(dt)\left(u_{n}(t)+\frac{A_{1n}}{6}\right)+I_{n}\frac{(A_{2n}+A_{3n})}{3}+\frac{A_{4n}}{3}\ ,$ (29)
@@
$\begin{cases}A_{1n}=dt(f_{n}+G_{n}[\bm{u}])\\ A_{2n}=dt(f_{n}+G_{n}[\bm{I}\circ(\bm{u}+\frac{\bm{A}_{1}}{2})])\\ A_{3n}=dt(f_{n}+G_{n}[\bm{I}\circ\bm{u}+\frac{\bm{A}_{2}}{2}])\\ A_{4n}=dt(f_{n}+G_{n}[\bm{I}\circ\bm{I}\circ\bm{u}+\bm{I}\circ\bm{A}_{3}])\end{cases}$ (30)
@@
where $m$ indicates the measured shell.
Appendix B Scale Aware Inflation Strategy
As a Monte Carlo-based method, EnKF is inherently prone to instability issues due to sampling errors. These arise from the inevitable limited number of ensemble members in comparison to what is needed to properly describe the system’s effective degrees of freedom [7]. For Gaussian and linear systems, the skill of stochastic EnKF monotonically converges to the ideal KF in the limit of an infinitely large ensemble. The minimum number of members required for satisfactory performance and computational affordability depends on the model-and- observational scenarios, as well as on the type of EnKF. For chaotic dynamics, in principle, there is a relation with the number of unstable modes of dynamics [9]. When the ensemble size is small, the estimated covariance matrix tends to contain spurious correlations, typically manifesting as underestimated variances and overestimated cross-correlations at large distance. While the latter are usually understood in physical space, the same is often the case also in spectral domain, whereby the actual correlation between scales/frequencies decreases with the frequency gap, while the ensemble-based correlations do not decrease similarly fast or not at all. These artificial, non-physical correlations pose a serious problem: when a new observation is assimilated, the update step may inadvertently modify variables that should remain unaffected, excessively reducing their associated error variance. The EnKF assigns insufficient weight to incoming observations, leading to filter divergence, which is particularly problematic in chaotic systems [48]. The resulting error perturbations increase the stiffness of the system and can ultimately cause catastrophic divergence, a finite-time numerical blow-up of the forecast [17].
The indisputable success of EnKF methods in high-dimension chaotic systems is also related to the development of a plethora of approaches, and ad-hoc fixes, to counteract the effect of the sampling error. These efforts can all be framed under two algorithmic solutions: inflation and localization [7, 8, 16]. Localization consists in restricting the effect of the analysis update to the vicinity (or local domain) of each observation and is a widely used and often indispensable technique in medium- to large-scale systems. The size of the localization domain is usually chosen based on practical arguments and may require numerical tuning, although recent studies have introduced scale-dependent localization approaches [49, 50].
In the shell model, the energy cascade is inherently local, with interactions occurring primarily between neighboring scales. This makes it natural to impose a tridiagonal structure on the covariance matrix as a form of spectral localization. However, our experiments reveal that such localization—either in the form of a tridiagonal approximation or of a stricter spectral correlation cutoff—while successful in eliminating spurious correlations due to sampling error, leads to physically imbalanced analyses causing larger forecast errors. We therefore omit using localization and instead shall demonstrate that an appropriate inflation strategy will suffice. We introduce a novel scale-aware inflation approach that mitigates the effects of undersampling without sacrificing physical consistency.
Inflation consists of artificially increasing the ensemble-based error covariance [19], and is usually performed in two alternative ways, additive or multiplicative [7]. It can be applied to either the forecast (prior) or the analysis (posterior) error covariance, with different effects.
(a) Measurement Sparseness. 6, 7, 8 $\lambda_{i}$ $\Delta t_{\text{obs}}=\tau_{15}=0.002\tau_{0}$ $\lambda_{14}=0.2$ $\Delta t_{\text{obs}}=\tau_{9}=0.02\tau_{0}$ $\lambda_{3}=0.1$ $\Delta t_{\text{obs}}=\tau_{8}=0.04\tau_{0}$ $\lambda_{5}=\lambda_{7}=\lambda_{11}=\lambda_{15}=\lambda_{16}=0.05$ $\lambda_{10}=0.01$ $\Delta t_{\text{obs}}=\tau_{7}=0.06\tau_{0}$ $\lambda_{1}=\lambda_{8}=\lambda_{12}=\lambda_{14}=\lambda_{15}=0.01$ $\lambda_{5}=\lambda_{11}=0.05$ $\lambda_{13}=0.005$ $\Delta t_{\text{obs}}=\tau_{6}=0.1\tau_{0}$ $\lambda_{i}=0$ $\Delta t_{\text{obs}}=\tau_{4}=0.2\tau_{0}$ $\lambda_{i}=0$
(b) Measured Scale Schemes. $\Delta t_{\text{obs}}=\tau_{15}=0.002\tau_{0}$ $\lambda_{i}$ 6, 14 $\lambda_{i}=0$ 4, 14 $\lambda_{i}=0$ 2, 14 $\lambda_{i}=0$ 6 $\lambda_{i}=0$ 6, 7 $\lambda_{2}=0.1$ $\lambda_{3}\!=\!\lambda_{4}\!=\!\lambda_{5}\!=\!\lambda_{7}\!=\!\lambda_{9}\!=\!\lambda_{13}\!=\!\lambda_{14}\!=\!0.25$ 6, 7, 8 $\lambda_{14}=0.2$ 6, 9, 10 $\lambda_{5}=\lambda_{10}=0.1$ 6, 11, 12 $\lambda_{5}=\lambda_{7}=0.2$, $\lambda_{14}=0.25$
Table 3: Inflation strength coefficients $\lambda_{i}$ used in both experiments: (a) varying observation sparseness $\Delta t_{\text{obs}}$ while measuring shells 6, 7, 8; (b) fixing $\Delta t_{\text{obs}}$ and varying the set of measured shells. Cases where $\lambda_{i}=0$ indicate that no inflation was required.
In additive inflation, a random term is added to each ensemble member to maintain the diversity and the spread of the ensemble. In multiplicative inflation, the ensemble-based error covariance, either prior or posterior, is inflated by a scalar factor. This inflates the ensemble variance by rescaling each posterior ensemble member around the mean as:
$g_{n}(t)\tilde{U}_{n}+(1-g_{n}(t))\tilde{\mu}_{n}\to\tilde{U}_{n}\ ,$ (32)
where $\tilde{\mu}_{n}$ is the posterior mean and $g_{n}(t)\geq 1$ is the inflation coefficient. This transformation increases the dispersion of the ensemble around the mean and rescales the covariance element $\tilde{\Sigma}_{nm}$ to $g_{n}g_{m}\tilde{\Sigma}_{nm}$.
Sacher and Bartello [18] showed that the sampling error in the Kalman gain estimate tends to scale directly proportional to the magnitude of the gain itself. This implies that larger corrections from observations require stronger inflation to maintain stability. It is therefore desirable to inflate the ensemble more in regions where observations exert a stronger influence. Motivated by this, we propose a scale-aware inflation scheme, where the spread reduction is compensated by:
with $\hat{\Sigma}_{nn}$ and $\tilde{\Sigma}_{nn}$ denoting the prior and posterior variances, respectively, and $\lambda\in{\mathbb{R}}^{+}$ the inflation strength, which is a hyperparameter to be tuned. This formulation is inspired by the original work of Zhang et al. [51], which used a similar expression but normalized by the posterior variance $\tilde{\Sigma}_{nn}$ instead of ${\hat{\Sigma}_{nn}}$. Since, statistically, $\langle\hat{\Sigma}_{nn}\rangle_{T}\geq\langle\tilde{\Sigma}_{nn}\rangle_{T}$, our version yields a smaller inflation coefficient for the same spread variation and $\lambda$, resulting in a less aggressive inflation.
The choice of the parameter $\lambda$ is dependent on the system and must be carefully tuned to ensure effective assimilation. Although our inflation strategy is scale-sensitive, tuning $\lambda$ adds computational cost. A detailed analysis of its effect on analysis statistics, filter stability, and comparison with other techniques lies beyond the scope of this work. However, even with extensive adjustment of $\lambda$, the proposed inflation strategy turns out to be very successful. It has been effective in stabilizing the assimilation process and in substantially reducing the minimum ensemble size needed to achieve satisfactorily performance and avoid filter divergence. Further and dedicated developments of the proposed scale-aware inflation method will be the content of follow-up works.

### v2

@@ VII Comparison with En4D-Var
Due to the chaotic nature of the shell model, any error - whether from the initial guess or from linearization - grows exponentially over time. As a result, the length of the assimilation window $T_{a}$ is typically limited to the eddy turnover time of the fastest measured shell. This poses a challenge for assimilating slower (i.e., lower-wavenumber) shells, which require longer windows. To address this, the 4D-Var process is performed sequentially: a long assimilation interval is divided into smaller windows, and background information is passed from one window to the next. A widely used strategy in numerical weather prediction to perform such sequential assimilation is the ensemble-based 4D-Var (En4D-Var) [25]. In this approach, multiple 4D-Var optimizations $\tilde{\bm{u}}^{(j)}(t)$ (with $j=1,\dots,L$ denoting the ensemble index) are run in parallel. Each member uses the background information $\hat{\bm{u}}^{(j)}(0)$, which comes from the posterior solution of the previous sequence, and randomly perturbed observations, the random perturbations are drawn from the same Gaussian distribution as the observation error itself. The background-error covariance matrix $\hat{\bm{\Sigma}}_{\mathbb{C}}=\langle(\hat{\bm{u}}-\langle\hat{\bm{u}}\rangle_{L})(\hat{\bm{u}}-\langle\hat{\bm{u}}\rangle_{L})^{\dagger}\rangle_{L}$ propagates the flow-dependent information across assimilation cycles. The modified cost function used in this ensemble-sequential version is:
$aligned J^{(j)}_{n_{\text{seq}}}=&\left\|\hat{\bm{u}}^{(j)}(n_{\text{seq}}T_{a})-\tilde{\bm{u}}^{(j)}(n_{\text{seq}}T_{a})\right\|_{\hat{\bm{\Sigma}}_{\mathbb{C}}}\\ +\sum_{t=n_{\text{seq}}T_{a}}^{(n_{\text{seq}}+1)T_{a}}&\left\|\bm{z}(t)-\bm{H}_{c}\tilde{\bm{u}}^{(j)}(t)-\bm{v}^{(j)}(t)\right\|_{\bm{R}_{\mathbb{C}}}\,$, (24)
@@
For the first sequence ($n_{\text{seq}}=0$), an ensemble $\hat{\bm{u}}^{(j)}$ is initialized as in Eq. (25) below and 4D-Var calculations are solved for a maximum of 50 iterations or until the reduction in the cost function reaches a tolerance. For later sequences ($n_{seq}>0$), the initial guess is the same as the background information, which is available from the previous sequence as $\hat{\bm{u}}^{(j)}(n_{seq}T_{a})=\tilde{\bm{u}}^{(j)}(n_{seq}T_{a})$. We also note that the first sequence does not have background information provided.
@@
$\hat{u}_{n}^{(j)}(t_{0})=\cases{\sqrt{\hfil}}{\langle|u_{n}|^{2}\rangle_{T}}\left[\cos(\phi_{n}^{(j)})+i\sin(\phi_{n}^{(j)})\right]&n\neq m\\ u_{n}(t_{0})+\epsilon_{n}^{(j)}&n=m\ \ ,$ (25)
@@
Regarding the execution of the algorithm, a regularization matrix $\mathcal{P}$ is added to the Hessian, where $\mathcal{P}$ is a diagonal matrix defined as $\mathcal{P}(n,n)=N_{\text{reg}}$ for $n<15$ and $\mathcal{P}(n,n)=N_{\text{reg}}/(2E_{n})$ otherwise. This matrix does not contribute to the cost function or its gradient, but it helps to regularize the Hessian inversion.
@@
The EnKF is run with $L=1000$ ensemble members, while En4D-Var uses $L=40$ for the $n=1,2,3$ case, and L=160L=160320320 for the $n=6,7,8$ case. The assimilation window for each sequence is $T_{a}=1000\,dt\approx 0.02\,\tau_{0}\approx\tau_{9}$. Similarly to EnKF, inflation could also be used in En4D-Var calculations. This may help to achieve the predictions with fewer ensemble members, but our results with inflation were inconclusive. This is because inflation is empirical and requires fine-tuning the inflation parameter, which is computationally expensive for En4D-Var calculations. Therefore, we do not use inflation for En4D-Var calculations.
@@
In summary, ensemble-based methods are necessary to assimilate scales beyond those directly measured. Both EnKF and En4D-Var perform well when three consecutive shells are observed. However, the current implementation of En4D-Var is significantly more computationally expensive than the stochastic EnKF method, which is parallelisable in both the forecast and ensemble steps. A more practical implementation of En4D-Var could employ a low-resolution model and quasi-Newton optimization in the inner loop. Such a study on trade-off between accuracy and computational cost is beyond the scope of the present study. Furthermore, 4D-Var also requires constructing the adjoint model, which can be challenging in complex flow systems.
@@ VIII Conclusions
Another promising direction involves integrating inflation optimization into a machine learning training framework. Since inflation effectively mimics the effect of larger ensemble sizes, learning the optimal inflation coefficients would allow us to simulate the behavior of massive ensembles without incurring their computational cost. A possible strategy is to adopt a solver-in-the-loop approach [53, 54], where inflation parameters are updated through backpropagation throughout the assimilation time window, rather than in a single analysis step. This would enable a global optimization of the assimilation process and open new pathways for hybrid dynamical–ML data assimilation methods [55].
(a) Measurement Sparseness.
6, 7, 8 $\lambda$ $N_{\text{inflation}}$
$\Delta t_{\text{obs}}=\tau_{15}=0.002\tau_{0}$ $\lambda=0.2$ 11
$\Delta t_{\text{obs}}=\tau_{9}=0.02\tau_{0}$ $\lambda=0.1$ 11
$\Delta t_{\text{obs}}=\tau_{8}=0.04\tau_{0}$ $0.01\leq\lambda\leq 0.05$ 66
$\Delta t_{\text{obs}}=\tau_{7}=0.06\tau_{0}$ $0.005\leq\lambda\leq 0.05$ 88
$\Delta t_{\text{obs}}=\tau_{6}=0.1\tau_{0}$ $\lambda=0$ 0
$\Delta t_{\text{obs}}=\tau_{4}=0.2\tau_{0}$ $\lambda=0$ 0
(b) Measured Scale Schemes.
$\Delta t_{\text{obs}}=\tau_{15}=0.002\tau_{0}$ $\lambda$ $N_{\text{inflation}}$
6 $\lambda=0$ 0
6, 7 $0.1\leq\lambda\leq 0.25$ 88
6, 7, 8 $\lambda=0.2$ 11
6, 9, 10 $\lambda=0.1$ 22
6, 11, 12 $0.2\leq\lambda\leq 0.25$ 33
Table 3: Range of inflation strength coefficients $\lambda$ used in both experiments:(a) varying the observation sparseness $\Delta t_{\text{obs}}$ while measuring shells 6, 7, and 8;(b) fixing $\Delta t_{\text{obs}}$ and varying the set of measured shells. In the last column number of experiments (out of 16) that need for inflation ($\lambda=0$ indicates that no inflation is required).
@@ Appendix A Runge-Kutta with exponential factor
$u_{n}(t+dt)=I_{n}^{2}(dt)\left(u_{n}(t)+\frac{A_{1n}}{6}\right)+I_{n}\frac{(A_{2n}+A_{3n})}{3}+\frac{A_{4n}}{6}\ ,$ (31)
@@
$\cases{A}_{1n}=dt(f_{n}+G_{n}[\bm{u}])\\ A_{2n}=dt(f_{n}+G_{n}[\bm{I}\circ(\bm{u}+\frac{\bm{A}_{1}}{2})])\\ A_{3n}=dt(f_{n}+G_{n}[\bm{I}\circ\bm{u}+\frac{\bm{A}_{2}}{2}])\\ A_{4n}=dt(f_{n}+G_{n}[\bm{I}\circ\bm{I}\circ\bm{u}+\bm{I}\circ\bm{A}_{3}])$ (32)
@@
where $m$ indicates the measured shell. The state is evolved by using the same as in Eq. (31) but with coefficient
$\cases{A}_{1n}=dt(f_{n}+G_{n}[\bm{u}]+T_{n}(\bm{z},t))\\ A_{2n}=dt(f_{n}+G_{n}[\bm{I}\circ(\bm{u}+\frac{\bm{A}_{1}}{2})]+T_{n}(\bm{z},t+\frac{dt}{2}))\\ A_{3n}=dt(f_{n}+G_{n}[\bm{I}\circ\bm{u}+\frac{\bm{A}_{2}}{2}]+T_{n}(\bm{z},t+\frac{dt}{2}))\\ A_{4n}=dt(f_{n}+G_{n}[\bm{I}\circ\bm{I}\circ\bm{u}+\bm{I}\circ\bm{A}_{3}]+T_{n}(\bm{z},t+dt))$ (34)
where
$aligned\bm{I}=\{&e^{-dt(\nu k_{0}^{2}+\delta_{0m}\alpha_{0})/2},e^{-dt(\nu k_{1}^{2}+\delta_{1m}\alpha_{1})/2},\ldots,\\ &e^{-dt(\nu k_{N-1}^{2}+\delta_{N-1,m}\alpha_{N-1})/2}\}\ .$
Appendix B Numerical Details on Inflation